MSG Corporation has $1,000,000 of 10-year, 6% bonds outstanding on December 31, 2018. The bonds have 3 years remaining to maturity. Assume interest is paid at the end of each month. The unamortized premium remaining on these bonds was $60,000. MSG uses straight-line amortization, so the unamortized premium would be $40,000 on December 31, 2019, provided none of the bonds had been retired before that day. 2. On May 1, 2019, $100,000 of the bonds were retired at 112. How much, and what type of gain or loss, results from the retirement?

A) $6,667 nonoperating loss B) $6,667 operating loss C) $6,667 nonoperating gain D) $6,667 operating gain
Answer #1

Answer:

B) $6,667 operating loss

Calculation:

Here, first we need to calculating the gain or loss. For that we need to find the payment at redemption and deduct the carrying value.

Paid at redemption is calculated by multiplying the amount at which bonds were retired.

Paid at redemption = 100,000 x 112% = 112,000

Since it was retired on May 1, we need to take 4 months out of 12 Months and also 1 year in 3 year maturity. So the carrying value will be:

Carrying Value = 1000,000 + 60,000 - (60,000 x 1/3 x 4/12) x 10% = 105,333

Operating Loss = Paid at redemption - Carrying Value = 112,000 - 105,333 = 6,667

Since it is the ordinary operations of a business, the loss incurred can be treated as operating loss.